# Kai Shu, Lu Cheng, Huan Liu

### v3

@@ 1 Introduction
While these automatic metrics are computationally efficient and used in many generation applications Zhang et al. (2022, 2023a, 2024c), their reliance on n-gram matching and reference-based designs significantly limits their applicability in dynamic and open-ended scenarios Liu et al. (2016); Reiter (2018). With the rise of deep learning models Devlin et al. (2019); Reimers and Gurevych (2019), many embedding-based assessment methods (e.g., BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021)) have also emerged. Although these small model-based metrics shift from word-level to embedding-level representations and offer greater flexibility, they still struggle to capture subtle attributes like helpfulness and harmlessness beyond relevance.
@@
Ye et al. (2024b), R-judge Yuan et al. (2024b), Do-not-answer Wang et al. (2024i), RAIN Li et al. , leaf, text width=41em ] ] [ Reliability
@@
 Wei et al. (2024b), Self-Alignment for Factuality Zhang et al. (2024g), FaithScore Jing et al. (2024), FENCE Xie et al. (2024b), leaf, text width=41em ] ] [ Relevance
@@
Thomas et al. (2024), Ma et al. (2024), LLMRank Hou et al. (2024), LLM Evaluation Chiang and Lee (2023), Yang and Lin (2024), Chen et al. (2024a)
@@
3.6) [ Gao et al. (2023), Just-Eval Lin et al. (2023), ICE Jain et al. (2023a), LLM-Eval Lin and Chen (2023b), GEMBA Kocmi and Federmann (2023),
KIEVAL Yu et al. (2024b), OAIF Guo et al. (2024), Comp-Analysis Zhang et al. (2024a), LostITS Huang et al. (2024), leaf, text width=41em ] ] ] [ Methodology
@@
PROMETHEUS2 Kim et al. (2024), InstructScore Xu et al. (2023), , leaf, text width=25em ] ] ] [ Tuning Techniques
@@
Meta-RewardingWu et al. (2024a), Self-Taught Wang et al. (2024f), PORTIALi et al. (2024n), leaf, text width=25em ] ] ] ] [ Prompting
@@
4.2.1) [ MT-Bench Zheng et al. (2023), RLAIF Lee et al. (2023), SALMON Sun et al. (2024), SELF-JUDGE Lee et al. (2024),
@@
Beigi et al. (2024a), AUTO-J Li et al. (2024e), Bai et al. (2023a), Gao et al. (2023), Prometheus Kim et al. , KIEVAL
Yu et al. (2024b), CEB Wang et al. (2024e), Murugadoss et al. (2024), Liu et al. (2024b), OAIF Guo et al. (2024),
SALMON Sun et al. (2024), SELF-JUDGE Lee et al. (2024), DALK Li et al. (2024c), Qian et al. (2024),
RevisEval Zhang et al. (2024f), LLM-as-a-personalized-judge Dong et al. (2024), Li et al. (2024k), Li et al. (2024h), leaf, text width=33em ] ] [ Multi-Agent
Collaboration (§4.2.3) [ PRD Li et al. (2023), Zhang et al. (2023b), Wu et al. (2023), MPA Zhu et al. (2024b), JudgeLM Zhu et al. (2023),
@@
Jung et al. (2024), Fellowship Arif et al. (2024), MATEval Li et al. (2024l), Zhang et al. (2024e), leaf, text width=33em ] ] [ Demonstration
@@
4.2.5) [ LLM-as-an-examine Bai et al. (2023b), KIEVAL Yu et al. (2024b), Auto-Arena Zhao et al. (2024c),  Moniri et al. (2024), leaf, text width=33em ] ] [ Comparison Acceleration
@@
5.1) [ Bi et al. (2023), Fei et al. (2023), Zhou et al. (2023), Wang et al. (2023a), Nan et al. (2024), Zheng et al. (2023), Gao et al. (2023), Wu et al. (2023),
@@
5.2) [ Bai et al. (2022), Lee et al. (2023), Sun et al. (2024), Guo et al. (2024), Arif et al. (2024), Li et al. (2024i), Yuan et al. (2024c), Wu et al. (2024a),
Pace et al. (2024), Lee et al. (2024), Tong et al. (2024), Zhai et al. (2024), Liu et al. (2024c), Liang et al. (2024c), Zhang et al. (2024g),
@@
Zhou et al. (2024b), Lahoti et al. (2023), Liang et al. (2023), Li et al. (2024b), Besta et al. (2024), Ong et al. (2024), Zhao et al. (2024b),
Kawabata and Sugawara (2024), Xie et al. (2024b), Lightman et al. (2023), Li et al. , Setlur et al. (2024), leaf, text width=41em ] ] ] ]
@@
The recent advanced large language models (LLMs) such as GPT-4 Achiam et al. (2023) and o1 Ope , have demonstrated striking performance in instruction following, query understanding and response generation. This progress has motivated researchers to propose the concept of “LLM-as-a-judge” Zheng et al. (2023), which leverages the powerful LLMs to score, rank and select from a group of candidates (Figure 1). The strong performance of LLMs Brown et al. (2020) combined with well-designed assessment pipelines Li et al. (2023); Beigi et al. (2024b); Bai et al. (2023a) leads to fine-grained and detailed judgment for various evaluation applications, significantly addressing the limitations of traditional evaluation methods, setting a new standard for what is achievable in NLP evaluation. Beyond evaluation, LLM-as-a-judge has also been widely adopted across the entire LLM lifecycle, including tasks like alignment Bai et al. (2022); Lee et al. (2023), retrieval Li and Qiu (2023); Li et al. (2024c), and reasoning Liang et al. (2023); Zhao et al. (2024b). It empowers LLMs with a series of advanced capabilities such as self-evolution Sun et al. (2024), active retrieval Li et al. (2024c) and decision-making Yang et al. (2023), driving their transformation from traditional models to intelligent agents Zhuge et al. (2024). However, as LLM-as-a-judge develops rapidly, challenges like judging bias and vulnerability Koo et al. (2023a); Park et al. (2024); Dai et al. (2024) are also emerging. A systematic review of current techniques and future challenges would therefore be valuable for advancing LLM-based judgment methods.
@@
Differences from Other LLM-related Surveys. LLMs have been a popular topic in recent years and got many related surveys Zhao et al. (2023b); Chang et al. (2024); Xiong et al. (2024a). While there are several surveys focusing on LLM-based evaluation for natural language generation (NLG) evaluation Gao et al. (2024a); Li et al. (2024m), this work aims to provide a comprehensive overview of the LLM-as-a-judge approach. As mentioned, LLM-as-a-judge has already been applied in a broader range of scenarios beyond evaluation, making it essential to offer a general perspective for its summarization and categorization. Besides, there are also some surveys focus on LLM-powered applications, such as LLMs-based data annotation Tan et al. (2024b), data augmentation Zhou et al. (2024c), and self-correction Pan et al. (2024). However, there is still a lack of a systematic and comprehensive survey dedicated to LLM-as-a-judge.
@@ 2.1 Input
When $n=1$, it becomes a point-wise judgment where the judge LLMs will solely focus on one candidate sample Gao et al. (2023).
@@ 2.2 Output
In ranking-based judgment, the output is a ranking of each candidate sample, represented as $R=\{C_{i}>...>C_{j}\}$. This comparative approach is useful in scenarios where establishing a rank order among candidates is required Li et al. (2023).
@@ 3.2 Harmlessness
Evaluating harmlessness in text data is important for both content moderation and for creating or curating synthetic datasets. Given how expensive and time-consuming human labeling efforts are, and inspired by prior work that looks into moral beliefs embedded in LLMs Scherrer et al. (2023), many recent works have investigated the use of LLMs to evaluate harmlessness. State-of-the-art LLMs are capable of being used effectively for content moderation, either off the shelf when guided with some policy guidelines [^3] https://openai.com/index/using-gpt-4-for-content-moderation/, or when fine-tuned on safe/unsafe data Inan et al. (2023); Zhang et al. (2024f) [^4] https://about.fb.com/news/2023/12/purple-llama-safe-responsible-ai-development/Ye et al. (2024b) explore the feasibility of using LLMs to evaluate harmlessness in a fine-grained manner, among other attributes, and find that proprietary models perform much better than open source ones. Wang et al. (2024i) use OpenAI’s GPT-4 to evaluate harmlessness and further compare the performance with that of a smaller pre-trained language model fine-tuned for this specific task. Additionally, Bai et al. (2022) use principles to guide LLMs to make harmlessness evaluations for alignment purposes, a paradigm they call Constitutional AI. Phute et al. (2023) use the same LLM to evaluate its responses as harmful, and provide insights and best practices for GPT-3.5 and Llama-2. Xie et al. (2024a) perform a comprehensive comparison of several LLMs on a benchmark of LLM safety refusals and find small LLMs are effective safety judges when used in fine-tuned settings. During inference time, Li et al. propose Rewindable Auto-regressive INference (RAIN), which allows LLMs to conduct self-evaluation and rewind for AI safety.
@@ 3.3 Reliability
Reliability is a crucial attribute for LLMs, enabling them to generate factual and faithful content while also expressing uncertainty or acknowledging gaps in knowledge about certain topics. Regarding factuality, Wang et al. (2024a) introduce HALU-J, a critique-based hallucination judge that enhances factuality assessment by selecting relevant evidence and providing detailed critiques.  Cheng et al. (2023) design an automated evaluation method using GPT-4 to judge whether a model’s output is hallucinated. Additionally, several works adopt judge LLMs for long-form factuality evaluation. In the context of dialogues, Luo et al. (2024) collect a large-scale benchmark for automatic dialogue-level hallucination evaluation. Based on this dataset, they introduce HalluJudge, a specialized judge language model for evaluating hallucinations at the dialogue level.  Min et al. (2023) develop FactScore, a fine-grained method for evaluating the factuality of long-form generation by first splitting content into atomic-level sentences and retrieving relevant corpus from Wikipedia to assess their factuality. Building on this approach, Wei et al. (2024b) propose equipping judge LLMs with a Google search API to enable more flexible and efficient factuality assessments. Jing et al. (2024) expand this fine-grained reliability evaluation to the multimodal area and introduce FaithScore. Zhang et al. (2024g) adopt a similar strategy in LLM alignment by creating a synthetic alignment dataset, which involves evaluating and filtering each generated sample using claim extraction and self-judging techniques. Xie et al. (2024b) train an external critique-based LLM as the judge to provide claim-level claim-level factuality feedback in the generation stage to improve response factuality. For uncertainty judgment, Xu et al. (2024c) present SaySelf, a novel training framework that teaches LLMs to express more fine-grained confidence estimates with self-consistency prompting and group-based calibration training.
@@ 3.4 Relevance
Relevance measures the extent to which generated or retrieved content aligns with the original query. Traditional relevance assessment methods often rely on keyword matching Robertson et al. (2009) or semantic similarity Gao et al. (2021), which struggle to capture subtle differences or nuances in context. Relevance assessment with judge LLMs has been explored and validated to be a more fine-grained and effective manner across various applications Chiang and Lee (2023). In conversation evaluation, Lin and Chen (2023a) first propose to replace the expensive and time-consuming human annotation with LLM judgment in relevance assessment, providing conversation context and generated response for the judge LLM to evaluate. Similarly, Abbasiantaeb et al. (2024) apply LLM-as-a-judge in conversation search, collaborating with human annotators to address issues related to incomplete relevance judgments. In retrieval-augmented generation (RAG) scenarios, Li and Qiu (2023) utilize LLMs to determine which historical memory is most relevant for solving the current problem. Following this concept, Li et al. (2024c) also propose to adopt LLM as a re-ranker to judge and filter out noisy and irrelevant knowledge in a sub-knowledge graph. Recently, LLM-as-a-judge has also been used in multimodal applications for relevance judgment Lin et al. (2025). Both Yang and Lin (2024) and Chen et al. (2024a) adopt a multimodal LLM to build the automatic assessment benchmark. In contrast, Chen et al. (2024e) employ the multimodal reward model to assess the relevance in text-to-image generation. In SQL generation evaluation, Zhao et al. (2023a) propose to leverage LLM to determine SQL equivalence. Besides, LLM-as-a-judge has also been explored in many traditional retrieval applications for relevance assessment, such as search Thomas et al. (2024), retrieval Ma et al. (2024) and recommendation Hou et al. (2024).
@@ 3.6 Overall Quality
As previously mentioned, LLM-as-a-judge can be employed to perform multi-aspect and fine-grained assessments across various tasks and applications. However, in many cases, a general assessment is still required to represent the candidates’ overall quality for comparison or ranking purposes. One straightforward approach to obtain this overall score is to calculate the average of the aspect-specific scores Lin et al. (2023); Lin and Chen (2023b). Additionally, several other studies present assessment results for each attribute and prompt the LLM judge to generate an overall quality judgement Yu et al. (2024b); Guo et al. (2024); Zhang et al. (2024a). Beyond summarizing overall quality from multiple attributes, other studies focus on directly producing an overall judgment. For instance, in traditional NLP tasks like summarization Gao et al. (2023); Jain et al. (2023a) and machine translation Huang et al. (2024); Kocmi and Federmann (2023), the evaluation dimensions are less diverse compared to more open-ended, long-form generation tasks. Consequently, LLM-as-a-judge is typically used in these cases to directly prompt the LLM to generate an overall judgment.
@@
PROME-THEUS2 Kim et al. (2024) Synthetic GPT-4 Preference Learning 300K SFT Joint Training, Weight Merging Mistral

### v4

@@ 1 Introduction
While these automatic metrics are computationally efficient and used in many generation applications Zhang et al. (2022, 2023a, 2024c), their reliance on n-gram matching and reference-based designs significantly limits their applicability in dynamic and open-ended scenarios Liu et al. (2016); Reiter (2018). With the rise of deep learning models Devlin et al. (2019); Reimers and Gurevych (2019), many embedding-based assessment methods (e.g., BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021)) have also emerged. Although these small model-based metrics shift from word-level to embedding-level representations and offer greater flexibility, they still struggle to capture subtle attributes Olson et al. (2024) like helpfulness and harmlessness beyond relevance.
@@
Ye et al. (2024b), R-judge Yuan et al. (2024b), Do-not-answer Wang et al. (2024i), RAIN lirain, leaf, text width=41em ] ] [ Reliability
@@
 Wei et al. (2024b), Self-Alignment for Factuality Zhang et al. (2024h), FaithScore Jing et al. (2024), FENCE Xie et al. (2024b), leaf, text width=41em ] ] [ Relevance
@@
Thomas et al. (2024), Ma et al. (2024a), LLMRank Hou et al. (2024), LLM Evaluation Chiang and Lee (2023), Yang and Lin (2024), Chen et al. (2024a)
@@
3.6) [ Gao et al. (2023b), Just-Eval Lin et al. (2023), ICE Jain et al. (2023a), LLM-Eval Lin and Chen (2023b), GEMBA Kocmi and Federmann (2023),
KIEVAL Yu et al. (2024c), OAIF Guo et al. (2024), Comp-Analysis Zhang et al. (2024a), LostITS Huang et al. (2024), leaf, text width=41em ] ] ] [ Methodology
@@
PROMETHEUS2 Kim et al. (2024b), InstructScore Xu et al. (2023), , leaf, text width=25em ] ] ] [ Tuning Techniques
@@
Meta-RewardingWu et al. (2024a), Self-Taught Wang et al. (2024f), PORTIALi et al. (2024o), leaf, text width=25em ] ] ] ] [ Prompting
@@
4.2.1) [ MT-Bench Zheng et al. (2023), RLAIF Lee et al. (2023), SALMON Sun et al. (2024a), SELF-JUDGE Lee et al. (2024),
@@
Beigi et al. (2024a), AUTO-J Li et al. (2024e), Bai et al. (2023a), Gao et al. (2023b), Prometheus Kim et al. , KIEVAL
Yu et al. (2024c), CEB Wang et al. (2024e), Murugadoss et al. (2024), Liu et al. (2024b), OAIF Guo et al. (2024),
SALMON Sun et al. (2024a), SELF-JUDGE Lee et al. (2024), DALK Li et al. (2024c), Qian et al. (2024),
RevisEval Zhang et al. (2024g), LLM-as-a-personalized-judge Dong et al. (2024), Li et al. (2024l), Li et al. (2024h), leaf, text width=33em ] ] [ Multi-Agent
Collaboration (§4.2.3) [ PRD Li et al. (2023a), Zhang et al. (2023b), Wu et al. (2023), MPA Zhu et al. (2024c), JudgeLM Zhu et al. (2023),
@@
Jung et al. (2024), Fellowship Arif et al. (2024), MATEval Li et al. (2024m), Zhang et al. (2024f), leaf, text width=33em ] ] [ Demonstration
@@
4.2.5) [ LLM-as-an-examine Bai et al. (2023b), KIEVAL Yu et al. (2024c), Auto-Arena Zhao et al. (2024c),  Moniri et al. (2024), leaf, text width=33em ] ] [ Comparison Acceleration
@@
5.1) [ Bi et al. (2023), Fei et al. (2023), Zhou et al. (2023), Wang et al. (2023a), Nan et al. (2024), Zheng et al. (2023), Gao et al. (2023b), Wu et al. (2023),
@@
5.2) [ Bai et al. (2022), Lee et al. (2023), Sun et al. (2024a), Guo et al. (2024), Arif et al. (2024), Li et al. (2024i), Yuan et al. (2024c), Wu et al. (2024a),
Pace et al. (2024), Lee et al. (2024), Tong et al. (2024), Zhai et al. (2024), Liu et al. (2024c), Liang et al. (2024c), Zhang et al. (2024h),
@@
Zhou et al. (2024c), Lahoti et al. (2023), Liang et al. (2023), Li et al. (2024b), Besta et al. (2024), Ong et al. (2024), Zhao et al. (2024b),
Kawabata and Sugawara (2024), Xie et al. (2024b), Lightman et al. (2023), lirain, Setlur et al. (2024), leaf, text width=41em ] ] ] ]
@@
The recent advanced large language models (LLMs) such as GPT-4 Achiam et al. (2023) and o1 Ope , have demonstrated striking performance in instruction following, query understanding and response generation. This progress has motivated researchers to propose the concept of “LLM-as-a-judge” Zheng et al. (2023), which leverages the powerful LLMs to score, rank and select from a group of candidates (Figure 1). The strong performance of LLMs Brown et al. (2020) combined with well-designed assessment pipelines Li et al. (2023a); Beigi et al. (2024b); Bai et al. (2023a) leads to fine-grained and detailed judgment for various evaluation applications, significantly addressing the limitations of traditional evaluation methods, setting a new standard for what is achievable in NLP evaluation. Beyond evaluation, LLM-as-a-judge has also been widely adopted across the entire LLM lifecycle, including tasks like alignment Bai et al. (2022); Lee et al. (2023), retrieval Li and Qiu (2023); Li et al. (2024c), and reasoning Liang et al. (2023); Zhao et al. (2024b). It empowers LLMs with a series of advanced capabilities such as self-evolution Sun et al. (2024a), active retrieval Li et al. (2024c) and decision-making Yang et al. (2023), driving their transformation from traditional models to intelligent agents Zhuge et al. (2024). However, as LLM-as-a-judge develops rapidly, challenges like judging bias and vulnerability Koo et al. (2023a); Park et al. (2024); Dai et al. (2024) are also emerging. A systematic review of current techniques and future challenges would therefore be valuable for advancing LLM-based judgment methods.
@@
Differences from Other LLM-related Surveys. LLMs have been a popular topic in recent years and got many related surveys Zhao et al. (2023b); Chang et al. (2024); Xiong et al. (2024a). While there are several surveys focusing on LLM-based evaluation for natural language generation (NLG) evaluation Gao et al. (2024a); Li et al. (2024n), this work aims to provide a comprehensive overview of the LLM-as-a-judge approach. As mentioned, LLM-as-a-judge has already been applied in a broader range of scenarios beyond evaluation, making it essential to offer a general perspective for its summarization and categorization. Besides, there are also some surveys focus on LLM-powered applications, such as LLMs-based data annotation Tan et al. (2024b), data augmentation Zhou et al. (2024d), and self-correction Pan et al. (2024). However, there is still a lack of a systematic and comprehensive survey dedicated to LLM-as-a-judge.
@@ 2.1 Input
When $n=1$, it becomes a point-wise judgment where the judge LLMs will solely focus on one candidate sample Gao et al. (2023b).
@@ 2.2 Output
In ranking-based judgment, the output is a ranking of each candidate sample, represented as $R=\{C_{i}>...>C_{j}\}$. This comparative approach is useful in scenarios where establishing a rank order among candidates is required Li et al. (2023a).
@@ 3.2 Harmlessness
Evaluating harmlessness in text data is important for both content moderation and for creating or curating synthetic datasets. Given how expensive and time-consuming human labeling efforts are, and inspired by prior work that looks into moral beliefs embedded in LLMs Scherrer et al. (2023), many recent works have investigated the use of LLMs to evaluate harmlessness. State-of-the-art LLMs are capable of being used effectively for content moderation, either off the shelf when guided with some policy guidelines [^3] https://openai.com/index/using-gpt-4-for-content-moderation/, or when fine-tuned on safe/unsafe data Inan et al. (2023); Zhang et al. (2024g) [^4] https://about.fb.com/news/2023/12/purple-llama-safe-responsible-ai-development/Ye et al. (2024b) explore the feasibility of using LLMs to evaluate harmlessness in a fine-grained manner, among other attributes, and find that proprietary models perform much better than open source ones. Wang et al. (2024i) use OpenAI’s GPT-4 to evaluate harmlessness and further compare the performance with that of a smaller pre-trained language model fine-tuned for this specific task. Additionally, Bai et al. (2022) use principles to guide LLMs to make harmlessness evaluations for alignment purposes, a paradigm they call Constitutional AI. Phute et al. (2023) use the same LLM to evaluate its responses as harmful, and provide insights and best practices for GPT-3.5 and Llama-2. Xie et al. (2024a) perform a comprehensive comparison of several LLMs on a benchmark of LLM safety refusals and find small LLMs are effective safety judges when used in fine-tuned settings. During inference time, Li et al. (2023b) propose Rewindable Auto-regressive INference (RAIN), which allows LLMs to conduct self-evaluation and rewind for AI safety.
@@ 3.3 Reliability
Reliability is a crucial attribute for LLMs, enabling them to generate factual and faithful content while also expressing uncertainty or acknowledging gaps in knowledge about certain topics. Regarding factuality, Wang et al. (2024a) introduce HALU-J, a critique-based hallucination judge that enhances factuality assessment by selecting relevant evidence and providing detailed critiques.  Cheng et al. (2023) design an automated evaluation method using GPT-4 to judge whether a model’s output is hallucinated. Additionally, several works adopt judge LLMs for long-form factuality evaluation. In the context of dialogues, Luo et al. (2024) collect a large-scale benchmark for automatic dialogue-level hallucination evaluation. Based on this dataset, they introduce HalluJudge, a specialized judge language model for evaluating hallucinations at the dialogue level.  Min et al. (2023) develop FactScore, a fine-grained method for evaluating the factuality of long-form generation by first splitting content into atomic-level sentences and retrieving relevant corpus from Wikipedia to assess their factuality. Building on this approach, Wei et al. (2024b) propose equipping judge LLMs with a Google search API to enable more flexible and efficient factuality assessments. Jing et al. (2024) expand this fine-grained reliability evaluation to the multimodal area and introduce FaithScore. Zhang et al. (2024h) adopt a similar strategy in LLM alignment by creating a synthetic alignment dataset, which involves evaluating and filtering each generated sample using claim extraction and self-judging techniques. Xie et al. (2024b) train an external critique-based LLM as the judge to provide claim-level claim-level factuality feedback in the generation stage to improve response factuality. For uncertainty judgment, Xu et al. (2024c) present SaySelf, a novel training framework that teaches LLMs to express more fine-grained confidence estimates with self-consistency prompting and group-based calibration training.
@@ 3.4 Relevance
Relevance measures the extent to which generated or retrieved content aligns with the original query. Traditional relevance assessment methods often rely on keyword matching Robertson et al. (2009) or semantic similarity Gao et al. (2021), which struggle to capture subtle differences or nuances in context. Relevance assessment with judge LLMs has been explored and validated to be a more fine-grained and effective manner across various applications Chiang and Lee (2023); Bach et al. (2024); Lu et al. (2024). In conversation evaluation, Lin and Chen (2023a) first propose to replace the expensive and time-consuming human annotation with LLM judgment in relevance assessment, providing conversation context and generated response for the judge LLM to evaluate. Similarly, Abbasiantaeb et al. (2024) apply LLM-as-a-judge in conversation search, collaborating with human annotators to address issues related to incomplete relevance judgments. In retrieval-augmented generation (RAG) scenarios, Li and Qiu (2023) utilize LLMs to determine which historical memory is most relevant for solving the current problem. Following this concept, Li et al. (2024c) also propose to adopt LLM as a re-ranker to judge and filter out noisy and irrelevant knowledge in a sub-knowledge graph. Recently, LLM-as-a-judge has also been used in multimodal applications for relevance judgment Lin et al. (2025); Zhang et al. (2024i); Zhu et al. (2024b); Tu et al. (2024); Ma et al. (2024b). Both Yang and Lin (2024) and Chen et al. (2024a) adopt a multimodal LLM to build the automatic assessment benchmark. In contrast, Chen et al. (2024e) employ the multimodal reward model to assess the relevance in text-to-image generation. In SQL generation evaluation, Zhao et al. (2023a) propose to leverage LLM to determine SQL equivalence. Besides, LLM-as-a-judge has also been explored in many traditional retrieval applications for relevance assessment, such as search Thomas et al. (2024), retrieval Ma et al. (2024a) and recommendation Hou et al. (2024).
@@ 3.6 Overall Quality
As previously mentioned, LLM-as-a-judge can be employed to perform multi-aspect and fine-grained assessments across various tasks and applications. However, in many cases, a general assessment is still required to represent the candidates’ overall quality for comparison or ranking purposes. One straightforward approach to obtain this overall score is to calculate the average or weighted average of the aspect-specific scores Lin et al. (2023); Lin and Chen (2023b); Saad-Falcon et al. (2024). Additionally, several other studies present assessment results for each attribute and prompt the LLM judge to generate an overall quality judgement Yu et al. (2024c); Guo et al. (2024); Zhang et al. (2024a). Beyond summarizing overall quality from multiple attributes, other studies focus on directly producing an overall judgment. For instance, in traditional NLP tasks like summarization Gao et al. (2023b); Jain et al. (2023a) and machine translation Huang et al. (2024); Kocmi and Federmann (2023), the evaluation dimensions are less diverse compared to more open-ended, long-form generation tasks. Consequently, LLM-as-a-judge is typically used in these cases to directly prompt the LLM to generate an overall judgment.
@@
PROME-THEUS2 Kim et al. (2024b) Synthetic GPT-4 Preference Learning 300K SFT Joint Training, Weight Merging Mistral